# Closed-Form Interpretation of Neural Network Latent Spaces with Symbolic Gradients

### v3

@@ 2 Evolve $T$ with (
3 aaa $g_{T}\leftarrow[\nabla T(\mathbf{x})\text{ for }\mathbf{x}\text{ in }X]$triangleright Gradients of symbolic model
@@ 4 Experiments
1 $2\times 2$ 4 Similarity Transformation Trace $A_{11}+A_{22}$ $\frac{A_{11}+A_{22}}{-0.878}$
2 Determinant $A_{11}A_{22}-A_{12}A_{21}$ $A_{12}A_{21}-A_{11}A_{22}$
3 $3\times 3$ 9 Similarity Transformation Trace $A_{11}+A_{22}+A_{33}$ $A_{11}+A_{22}+A_{33}$
4 $3\times 3$ Antisymmetric Sum of Principal Minors $A_{12}^{2}+A_{23}^{2}+A_{13}^{2}$ $A_{12}A_{21}+A_{23}A_{32}+A_{13}A_{31}$
5 $4\times 4$ 16 Similarity Transformation Trace $A_{11}+A_{22}+A_{33}+A_{44}$ $A_{11}+A_{22}+A_{33}+A_{44}$
6 6 Lorentz Transformation Inner Product $E_{1}B_{1}+E_{2}B_{2}+E_{3}B_{3}$ $E_{1}B_{1}+E_{2}B_{2}+E_{3}B_{3}$
@@ 3 $3\times 3$ 9 Similarity Transformation Trace $A_{11}+A_{22}+A_{33}$ $A_{11}+A_{22}+A_{33}$
7 1 $\frac{1}{2}x^{2}$ Energy $\frac{1}{2}v^{2}+\frac{1}{2}x^{2}$ $v^{2}+x^{2}$
8 $\sin(x)$ $\frac{1}{2}v^{2}+\sin(x)$ $v^{2}+\sin(x)+\sin(x)$
9 $\frac{1}{2}x^{2}+\frac{1}{4}x^{4}$ $\frac{1}{2}v^{2}+\frac{1}{2}x^{2}+\frac{1}{4}x^{4}$ $v\cdot v+x\cdot x+0.513(x\cdot x)^{2}$
10 $\frac{1}{2}x^{2}+\exp(x+1)$ $\frac{1}{2}v^{2}+\frac{1}{2}x^{2}+\exp(x+1)$ $\text{square}(v)+x\cdot x+\exp(x+1.684)$
11 2 $-r^{-2}$ Angular Momentum $x_{1}v_{2}-x_{2}v_{1}$ $x_{2}v_{1}-x_{1}v_{2}$
@@ 8 $\sin(x)$ $\frac{1}{2}v^{2}+\sin(x)$ $v^{2}+\sin(x)+\sin(x)$
12 4 Lorentz Transformation Spacetime Interval $t^{2}-x_{1}^{2}-x_{2}^{2}-x_{3}^{2}$ $x_{1}\cdot x_{1}+\text{square}(x_{2})-((t\cdot t)-\text{square}(x_{3}))$
We retrieve the invariants of matrices and various physical systems using our method. We consider invariants of matrices under similarity and Lorentz transformations. Additionally, we investigate dynamical systems characterized by a variety of potentials, as well as the invariants in Minkowski spacetime.
4.1 Datasets and Training
4.1.1 Invariants Under the Similarity Transformation
In experiments 1-3, 5 in Table 1, we search for the trace and determinant of matrices under the similarity transformation. Each data point is a triplet consisting of three matrices of dimension $n$: an anchor matrix $A$, a positive example $P$, and a negative example $N$. The anchor is sampled by generating a random matrix.
4.1.2 Invariants of Antisymmetric Matrices
For antisymmetric matrices in experiment 4, we prepare our dataset in the same way as we describe in 4.1.1. We first sample an antisymmetric $3\times 3$ matrix for the anchor $A$, followed by a similarity transformation for the positive sample $P$. Finally, we sample a new antisymmetric matrix for the negative sample $N$. While both the anchor and negative samples are antisymmetric, the positive sample does not inherit this property under the transformation $P=MAM^{-1}$ when $M$ is not orthonormal, because antisymmetry is not preserved under a general change of basis. Hence, we use all 9 entries of the matrix as input, although we acknowledge that one could easily enforce that $M$ is orthonormal, in which case only 3 inputs would be needed from each of $A$, $P$, and $N$.
We attempt to retrieve (8), which is invariant under the similarity transformation. Since we use the antisymmetric anchor matrix $A$ as input when computing $\nabla_{\mathbf{x}}f(\mathbf{x})$, we expect that the result of symbolic search would simplify to (10).
4.1.3 Invariants Under the Lorentz Transformation
In experiment 6, we apply the Lorentz transformation to the field strength tensor $F_{\mu\nu}$, which gives rise to the Lorentz invariants in Appendix A.1.3. Since the antisymmetry of $F_{\mu\nu}$ is preserved under the Lorentz transformation, each member of a triplet is antisymmetric, so we only use the 6 off-diagonal entries above (or equivalently below) the main diagonal as our input to the neural network. The anchor is a vector of these 6 entries from $F_{\mu\nu}$,
4.1.4 Potentials
The experiments in Table 2 correspond to motion in a potential, where we simulate trajectories by randomly sampling initial positions and velocities, and subsequently evolve these systems according to Hamilton’s equations. For each triplet $(x_{A},x_{P},x_{N})$, the anchor $x_{A}$ and positive sample $x_{P}$ are measurements at two different points along the same trajectory, while the negative sample $x_{N}$ is sampled from a different trajectory. The network must determine whether or not two measurements belong to the same particle.
4.1.5 Spacetime
In experiment 12 in Table 3, each triplet again consists of an anchor $x_{A}$, a positive sample $x_{P}$, and a negative sample $x_{N}$. The anchor is a randomly sampled four-vector representing an event in Minkowski spacetime. The positive sample is generated by applying a Lorentz transformation to the anchor, ensuring that the spacetime interval remains invariant. The negative sample, on the other hand, is another randomly generated four-vector that does not share the same spacetime interval as the anchor, allowing the neural network to distinguish between vectors that do and do not preserve this invariant.
@@ 4.2 Results
Figure 2: The latent space encodings of Siamese neural network applied to different data sets are compared with the corresponding ground truth concept for each data point. In all cases, it is possible to see a clear correlation. However, this correlation is mostly non-linear causing other interpretation methods to fail.
@@
We summarize the results of our experiments in Tables 13. For each experiment, we use the method outlined in section 3 to obtain a set of predicted expressions, which we present as a Pareto plot in Figure 4 and Figure 3. We identify the solution that most closely matches the correct expression, denoted as retrieved expression in tables 13. It is interesting to note that the correct expression is often the one corresponding to the steepest drop in the loss. We note that in some cases, the network learned a polynomial approximation to the desired expression, which we rectified by increasing the sampling range used to produce the dataset.
All obtained solutions match the correct expressions. It is possible for the symbolic search algorithm to instead return a solution that matches the ground truth one up to a piecewise invertible transformation, although we do not observe this in our experiments. We also note that the symbolic search algorithm may approximate the correct solution, or add simplifications to it. For example, the solution denoted by the striped green bar in Figure 4 (c) a uses $\exp(x_{1}\cdot x_{1})\approx 1+x^{2}+\frac{x^{4}}{2}$, which matches the correct solution up to the fourth order in $x$. In Figure 4 (d), the expression $2\exp(x+1)$ was simplified to $\exp(x+1+\ln(2))\approx exp(x+1.684)$.
@@
Furthermore, because the network’s latent layer consists of only a single neuron, we can directly compare the value it encodes for inputs $\mathbf{X}$ to the true underlying concept $g(\mathbf{X})$. We plot these quantities against each other in Figure 2. Note that these correlation plots are not a necessary component of our interpretation framework. We use them only to highlight the non-linear manner in which the neural network encodes the concept. In most experiments, the values encoded in the latent space are highly correlated with some well-known concept. In fact, the correlation plots for the trace are almost linear, which is expected as they can trivially be learned by a single-layer neural network with no non-linearities. In such cases, it may be possible to use other methods to interpret the neural network. However, most invariants are significantly more complex, and the neural network will encode them in a non-linear manner, in which case most other interpretation methods will fail.
Figure 4: The Pareto front for experiments involving matrices, summarizing the results of the symbolic gradient-based interpretation framework to find a candidate concept that is contained in the corresponding neural network latent space. Several possible equations are plotted in order of decreasing Mean Square Error (MSE) and increasing complexity. The red bar indicates the candidate that resembles the ground truth concept, which is often found at the point of steepest change of the Pareto front. The striped green bar denotes a solution that approximates the correct one up to the fourth order.
@@ 5 Conclusions
In this manuscript, we develop a framework to interpret any single neuron in neural network latent spaces in the form of a symbolic equation. It is based on employing symbolic search to find a symbolic tree that exhibits the same normalized gradients as the examined latent space neuron. We justify this procedure by defining an equivalence class of functions encoding the same concept, in which the membership criterion is that all members have parallel gradients at every point on the data manifold. Through this procedure, we enable the extraction of concepts encoded by latent space models.
We demonstrate the power of our approach by interpreting Siamese networks tasked with discovering invariants of matrices and conserved quantities of dynamical systems. We are able to uncover the correct equations in all of our examples. It is important to note that the symbolic search algorithm sometimes made clever approximations. For example, the anharmonic potential was summarized by an exponential function whose Taylor expansion agrees to fourth order in $x$. Further, the approach simplified expressions, for example, the term $2\exp(x+1)$ was transformed into $\exp(x+1+\ln(2))\approx\exp(x+1.684)$.
It is impossible to compare our results to other methods because our approach is the only general method that allows for the extraction of concepts encoded in latent spaces in closed form. As we have seen, sometimes the latent space encodings are approximately linearly correlated with the human-readable ground truth concept. In these cases, it is possible to retrieve the expression with traditional symbolic or polynomial regression. However, this is not the general case. It is important to note that there might exist publication bias towards linear encodings, since non-linear encodings cause traditional interpretations to fail.
@@
Appendix A Preliminaries
A.1 Matrix Invariants
A.1.1 Invariants of Rank-Two Tensors Under Similarity Transformations

### v4

@@ 2 Evolve $T$ with (
3 aaa $g_{T}\leftarrow[\nabla T(\mathbf{x})\text{ for }\mathbf{x}\text{ in }X]$ $\triangleright$ Gradients of symbolic model
@@
5 aaaaaaaaa $\textbf{ else }\nabla T(\mathbf{x})\text{ for }\nabla T(\mathbf{x})\text{ in }g_{T}$] $\triangleright$ Normalize Gradients
@@ 4 Experiments
4.1 Dataset Generation
To test the effectiveness of our method, we demonstrate it on 12 different datasets. Each dataset consists of $50000$ training triplets, $5000$ validation triplets and $10000$ test triplets. The triplets are constructed in the following way: once the anchor $x_{A}$ is sampled, the positive sample $x_{P}$ is obtained via $x_{P}=M(x_{A})$, where $M$ is a placeholder operator for a specific transformation that is defined for each experiment in Appendix E, and finally $x_{N}$ is sampled independently. The operation implemented by $M$ transforms $x_{A}$ to $x_{P}$ such that certain properties of $x_{A}$ are inherited by $x_{P}$, but the two points are otherwise unique. The invariants and conserved quantities corresponding to our data sets are: trace, determinant, sum of principal minors under the similarity transformation, the inner product and spacetime interval under the Lorentz transformation, and the energy and momentum in a variety of potentials. More details about each dataset, including how to reproduce them, can be found in Appendix E.
@@ 4.1 Dataset Generation
1 $2\times 2$ 4 Similarity Trace $A_{11}+A_{22}$ $0.677-A_{11}-A_{22}$
2 $2\times 2$ 4 Similarity Determinant $A_{11}A_{22}-A_{12}A_{21}$ $A_{12}A_{21}-A_{11}A_{22}$
3 $3\times 3$ 9 Similarity Trace $A_{11}+A_{22}+A_{33}$ $A_{11}+A_{22}+A_{33}$
4 $3\times 3$ Antisymmetric 9 Orthogonal Sum of Principal Minors $A_{12}^{2}+A_{23}^{2}+A_{13}^{2}$ $A_{12}A_{21}+A_{23}A_{32}+A_{13}A_{31}$
5 $4\times 4$ 16 Similarity Trace $A_{11}+A_{22}+A_{33}+A_{44}$ $A_{11}+A_{22}+A_{33}+A_{44}$
6 $4\times 4$ Field Strenth Tensor 6 Lorentz Determinant $E_{1}B_{1}+E_{2}B_{2}+E_{3}B_{3}$ $E_{1}B_{1}+E_{2}B_{2}+E_{3}B_{3}$
@@ 5 $4\times 4$ 16 Similarity Trace $A_{11}+A_{22}+A_{33}+A_{44}$ $A_{11}+A_{22}+A_{33}+A_{44}$
7 2 $\frac{1}{2}x^{2}$ Energy $\frac{1}{2}v^{2}+\frac{1}{2}x^{2}$ $0.173/(v^{2}+x^{2})$
8 2 $\frac{1}{2}x^{2}+\frac{1}{4}x^{4}$ Energy $\frac{1}{2}v^{2}+\frac{1}{2}x^{2}+\frac{1}{4}x^{4}$ $v^{2}+\exp(x^{2})$
9 2 $\sin(x)$ Energy $\frac{1}{2}v^{2}+\sin(x)$ $v^{2}+\sin(x)/(-0.5)$
10 2 $\frac{1}{2}x^{2}+\exp(x+1)$ Energy $\frac{1}{2}v^{2}+\frac{1}{2}x^{2}+\exp(x+1)$ $v^{2}+x^{2}+\exp(x)/\sin(0.184)$
11 4 $r^{-2}$ Angular Momentum $x_{1}v_{2}-x_{2}v_{1}$ $x_{2}v_{1}-x_{1}v_{2}$
@@
12 4 Lorentz Transformation Spacetime Interval $t^{2}-x_{1}^{2}-x_{2}^{2}-x_{3}^{2}$ $t^{2}-x_{1}^{2}-x_{2}^{2}-x_{3}^{2}$
@@ 4.2 Results
Figure 2: The latent space encodings of Siamese neural network applied to different data sets are compared with the corresponding ground truth concept for each data point. In all cases, it is possible to see a clear correlation. However, this correlation is mostly non-linear causing direct symbolic regression methods to fail, since they would attempt to fit additional variables for slopes and intercepts as well as the deformation to a non-linear dependency.
@@
For each experiment, we use the method outlined in section 3 to obtain a set of predicted expressions from the symbolic search algorithm, which we present as a Pareto chart in Figure 4 and Figure 3. The Pareto chart plots each of these expressions as a bar chart in decreasing order of loss and increasing complexity. Of these expressions, we identify the one that most closely matches the correct expression at the steepest drop of the Pareto front, translate the variables to the underlying experiments, and present it under the column titled retrieved expression in tables 13. The correct expression is typically found after the steepest drop in the loss, corresponding to the lowest complexity symbolic solution that captures the ground truth. A notable exception to this rule arises when the network learns a good approximation to the desired expression, which can be rectified by increasing the sampling range used to produce the dataset. Since $H_{g}$ contains many different symbolic solutions that are all connected by an invertible transformation, it is possible to find different formulations of the ground truth at different complexity levels.
In our experiments, it was possible to retrieve all the correct ground truth expressions or, in one example, a very good approximation. We also observe that the symbolic search algorithm might converge to a different formulation of the ground truth, by making simplifications to it or by expressing terms in a more complex manner.
Of note is experiment 7, where the symbolic search algorithm refuses to simplify the retrieved expression $0.173/(v^{2}+x^{2})$ to $v^{2}+x^{2}$. Further, in experiment 8, the resulting equation $v^{2}+\exp(x^{2})\approx v^{2}+1+x^{2}+\frac{x^{4}}{2}$ is an approximation which matches the correct solution up to the fourth order in $x$. Since we initialized the initial conditions for this experiment with a mean of $0.5$ the 6th order is suppressed by $1/3!\times 0.5^{6}\approx 0.0026$. We can confirm that by increasing the training data points to $200000$ the algorithm uncovers the correct ground truth. In experiment 10, $\sin(0.184)\approx 1/(2e)$ can be absorbed as prefactors and exponent leading to the correct discovery of the ground truth.
Figure 3: The Pareto front for experiments involving matrices, summarizing the results of the symbolic gradient-based interpretation framework to find a candidate concept that is contained in the corresponding neural network latent space. Several possible equations are plotted in order of decreasing Mean Square Error (MSE) and increasing complexity. The red bar indicates the candidate that resembles the ground truth concept, which is often found at the point of steepest change of the Pareto front.
We compare the latent projection $f(\mathbf{X})$ for all inputs on the data set $\mathbf{X}$ to the true underlying concept $g(\mathbf{X})$. This can be visualized by plotting these quantities against each other in Figure 2. Note that these correlation plots are not a necessary component of our interpretation framework and are solely used to highlight the non-linear manner in which the neural network encodes the concept. In most experiments, the values encoded in the latent spaces are highly correlated with the ground truth concepts. In fact, the correlation plots for the traces of the different matrices in Figure 2 are almost linear, which is expected as they can trivially be learned by a single-layer neural network with no non-linearities. In such cases, it is possible to use methods such as directly applying symbolic regression to the latent space to interpret the neural network. However, most invariants are significantly more complex, and the neural network will encode them in a non-linear manner, in which case interpretation by symbolic regression fails. All of these methods fail for the same reason - they attempt to retrieve the distorted version of the concept $\phi(g(\mathbf{x}))$, rather than the concept itself. In comparison, our method searches for a symbolic tree whose gradients are aligned with the network $f$. This means that the tree is not restricted to representing the distorted concept, and coupled with the complexity penalty of symbolic search, it yields the simplest possible expression whose gradients match the network $f$. We provide a comparison of our method to performing symbolic regression directly on the latent space Cranmer et al. (2020) in Appendix F, where only 7 of 12 experiments are successful. In our experiments, direct symbolic regression is capable of identifying invariants of polynomials up to second order. Of the 7 successful experiments, 3 of them are simply the trace, 2 are second order polynomials involving cross-terms, and the remaining two are also second order polynomials, but without any cross terms. Interestingly, direct symbolic regression manages to retrieve a valid version of the ground truth expression for the sum of principal minors of $3\times 3$ antisymmetric matrices, which is encoded in a highly non-linear manner according to the correlation plot in Figure 2. It is important to highlight that direct symbolic regression can rarely succeed in discovering the ground truth by accident even though it fails to approximate the latent projection. The reason behind this accidental success is likely based on finding a good solution on data-dense regions on the data manifold - see Figure 6 in Appendix F.
@@ 5 Conclusions
In this manuscript, we developed a framework to interpret any single neuron in neural network latent spaces in the form of a symbolic equation. It is based on employing symbolic search to find a symbolic tree that exhibits the same normalized gradients as the examined latent space neuron. The approach is suitable to interpret all kinds of neural networks applied to structured data within settings where concepts are formulated as scalar equations, like in science. The approach is limited by the expressibility of symbolic search algorithms and the challenge of isolating single neurons through bottlenecks or disentanglement.
We justify this procedure by defining an equivalence class of functions encoding the same concept, in which the membership criterion is that all members have parallel gradients at every point on the data manifold. Through this procedure, we enable the extraction of concepts encoded by latent space models.
We demonstrate the power of our approach by interpreting Siamese networks tasked with discovering invariants of matrices and conserved quantities of dynamical systems. We are able to uncover the correct equations or very good approximations in all of our examples. It is important to note that the symbolic search algorithm sometimes made clever approximations. For example, the anharmonic potential was summarized by an exponential function whose Taylor expansion agrees to fourth order in $x$.
It is impossible to compare our results to other methods because our approach is the only general method that allows for the extraction of concepts encoded in latent spaces in closed form. As we have seen, sometimes the latent space encodings are approximately linearly correlated with the human-readable ground truth concept. In these cases, it is possible to retrieve the expression with traditional symbolic or polynomial regression. However, this is not the general case. It is important to note that there might exist publication bias towards linear encodings, where direct symbolic or polynomial regression methods accidentally succeed.
@@
The code to create all data and perform the experiments in this work can be found at https://github.com/sjwetzel/PublicSymbolicGradientLatentSpaceInterpretation.
6 Acknowledgements
We thank the National Research Council of Canada for their partnership with Perimeter on the PIQuIL. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Economic Development, Job Creation and Trade. This work was supported by Mitacs through the Mitacs Accelerate program.
7 Contributions
SJW had the idea for the project and designed the research agenda. SJW provided the mathematical foundations, SJW and ZP were involved in writing code supporting this work. SJW and ZP wrote the manuscript and prepared figures.
Appendix A Appendix
Appendix B Equivalence of Equivalence Classes
B.1 Proof
Let $g(\mathbf{x}),f(\mathbf{x})\in C^{1}(D\subset\mathbb{R}^{n},\mathbb{R})$ be continuously differentiable functions ($C^{1}(D\subset\mathbb{R}^{n},\mathbb{R})$ is the vector space of differentiable functions from $D$ to $\mathbb{R}$) and $D\subset\mathbb{R}^{n}$ be the data manifold which is required to be compact and simply connected. Then $\tilde{H}_{g}=H_{g}=H_{g+}\cup H_{g-}$ where
$\tilde{H}_{g}=\left\{f(\mathbf{x})\in C^{1}(D\subset\mathbb{R}^{n},\mathbb{R})\ |\exists\text{ invertible }\phi\in C^{1}(\mathbb{R},\mathbb{R})\ :\ f(\mathbf{x})=\phi(g(\mathbf{x}))\right\}$ (7)
and
$H_{g\pm}=\left\{f(\mathbf{x})\in C^{1}(D\subset\mathbb{R}^{n},\mathbb{R})\ |\ \frac{\nabla f(\mathbf{x})}{\lVert\nabla f(\mathbf{x})\rVert}=\frac{\pm\nabla g(\mathbf{x})}{\lVert\nabla g(\mathbf{x})\rVert}\lor\nabla f(\mathbf{x})=\nabla g(\mathbf{x})=0,\ \forall\mathbf{x}\in D\right\}$ (8)
Proof: One can see that for each function $f\in\tilde{H}_{g}\ \phi:\nabla f(\mathbf{x})=\phi^{\prime}(g(\mathbf{x}))\nabla g(\mathbf{x})$ , hence the gradients are parallel and thus $\tilde{H}_{g}\subset H_{g}$.
It remains to be shown that for each function $f\in H_{g}\ \exists\phi:f(\mathbf{x})=\phi(g(\mathbf{x}))$. Let us focus on $f\in H_{g+}$, the proof is analogous for $H_{g-}$. Let us explicitly construct the function $\phi$ that maps between $f$ and $g$. Defining $\phi^{\prime}$ through
$\nabla f(\mathbf{x})=\phi^{\prime}(g(\mathbf{x}))\nabla g(\mathbf{x})$ (9)
omits the avoids the necessity of defining $\phi$ at locations where the gradients are zero. This definition leads to an integrable $\phi^{\prime}(g(\mathbf{x}))\nabla g(\mathbf{x})=\nabla f(\mathbf{x})$ because a) the images of $f(D)$ and $g(D)$ are compact, thus $\phi^{\prime}$ maps between compact subsets of $\mathbb{R}$ and b) $\phi^{\prime}$ is continuous. For any simply connected $D\subset\mathbb{R}^{n}$ we can define the $C^{1}$-curve $\mathbf{x}:[t_{0},t_{1}]\rightarrow D$, thus a variable transformation within the calculation of the contour integral yields:
$\phi(g(\mathbf{x}(t_{1})))-\phi(g(\mathbf{x}(t_{0})))$ $=\int_{g(\mathbf{x}(t_{0}))}^{g(\mathbf{x}(t_{1}))}\phi^{\prime}(g)\ dg$ (10)
$=\int_{\mathbf{x}(t_{0})}^{\mathbf{x}(t_{1})}\phi^{\prime}(g(\mathbf{x}))\nabla g(\mathbf{x})\cdot d\mathbf{x}$ (11)
$=\int_{t_{0}}^{t_{1}}\phi^{\prime}(g(\mathbf{x}(t)))\nabla g(\mathbf{x}(t))\cdot\dot{\mathbf{x}}(t)\ dt$ (12)
[…]
$=\int_{x(t_{0})}^{x(t_{1})}\nabla f(\mathbf{x})\ d\mathbf{x}$ (14)
$=\int_{f(x(t_{0}))}^{f(x(t_{1}))}\ df$ (15)
$=f(\mathbf{x}(t_{1}))-f(\mathbf{x}(t_{0}))$ (16)
Similarly, one can proof the existence of $\tilde{\phi}:\tilde{\phi}(f(\mathbf{x}))=g(\mathbf{x})$ such that $f(\mathbf{x})=\phi(\tilde{\phi}(f(\mathbf{x})))$ and thus $\phi$ is invertible. While this proof assumes $f\in H_{g+}$ it is analogously possible to construct $\phi$ for $f\in H_{g-}$. Having explicitly constructed $\phi$ proofs $H_{g}=H_{g+}\cup H_{g-}=\tilde{H}_{g}$.
B.1.1 Assumptions
In a practical machine learning applications not all assumptions from the prior section that ensure $H_{g}=\tilde{H}_{g}$ hold true. However, even then $H_{g}\approx\tilde{H}_{g}$ can provide a good approximation that allows for a retrieval of the function that a neuron encodes.
A machine learning data set approximate the data manifold $D\subset\mathbb{R}^{n}$. If there is a divergence in the function that the machine learning model is supposed to approximate, the data set might not be closed and thus not compact. A data manifold $D$ might not be simply connected, especially if it is in the form of categorical data or images.
A neural network classifier, if successfully trained, tends to approximate a categorical output, which is neither continuous nor differentiable. However, this binary output is typically an approximation mediated by sigmoid or softmax activation functions, which indeed are continuously differentiable. Still, interpreting artificial neural networks with the framework introduced in this paper experiences numerical artifacts if a gradient is taken from a network that contains sigmoid or softmax activation functions. For this reason, I suggest avoiding these activation functions in the design of hidden layers and removing them from the output neuron during the interpretation process (the same argument holds true for tanh or related activation functions).
The above definitions of equivalence classes could be extended to piecewise $C^{1}(\mathbb{R}^{n},\mathbb{R})$ functions. This function set contains many artificial neural networks that include piecewise differentiable activation functions like $\text{ReLU}(x)=\max(0,x)$. However, this causes problems when evaluating derivatives close to $\text{ReLU }(x)=0$. In practice, one can observe that piecewise $C^{1}$ activation functions lead to computational artifacts when calculating gradients. Hence, I suggest using $\text{ELU}=\exp(x)-1|x\leq 0,x|x>0$ as the preferred activation function in hidden layers.
Appendix C Preliminaries
C.1 Matrix Invariants
C.1.1 Invariants of Rank-Two Tensors Under Similarity Transformations